OpenStudy (anonymous):

e^x-squareroot e^x=6 please help, i also need e^2x-squareroot e^2x=6 but i imagine since theyre so similar, theyre solved the same way

OpenStudy (amistre64):

let e^x = r^2 for a substitution ... and see if it looks familiar afterwards

OpenStudy (amistre64):

r^2 -sqrt(r^2) = 6 r^2 -r = 6 is a quadratic, solve for r and then solve for r^2 = e^x

OpenStudy (amistre64):

same basic prinicple applies to the other one if you dont see it as a quad yet